If f = (1, 2), (2, 3), (3, 4), (4, 5),
g = (1, -2), (3, -3), (5, -5), and
h = (1, 0), (2, 1), (3, 2),

find the following and state the domain:

f - g
1
Expert's answer
2013-01-14T11:01:14-0500
Determine value of x where both function (f, g) are defined:
x = 1 -> f(1) = 2, g(1) = -2
x = 3 -> f(3) = 4, g(3) = -3 &

Find function f - g when x = 1, 3:

f - g(1) = f(1) - g(1) = 2 - (-2) = 4
f - g(3) = f(3) - g(3) = 4 - (-3) = 7

So f*g*h& = (1, 4), (3, 7)

domain of a function is the set of "input" or argument values (x) for which the function is defined, so

domain: {1, 3}